# Poisson Midpoint Method for Log Concave Sampling: Beyond the Strong Error Lower Bounds

### v2

@@ 4 Proof overview
    $||V||\leq\beta$ a.s., $\mathbb{E}[V]=0$, and $\mathbb{E}[VV^{T}]=\Sigma.$
@@
$W_{2}^{2}(X_{tK+i+1},\tilde{X}_{tK+i+1})\leq(1-\tfrac{\alpha\eta}{2K})W_{2}^{2}(\tilde{X}_{tK+i},X_{tK+i})+\tfrac{\eta^{4}L^{2}}{K}\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}+\tfrac{\eta^{4}L^{4}d}{\alpha K}$
@@
The proof in this case is similar to the overdamped case. Here we use the fact we can obtain a contraction map similar to $T$ in Equation (4) even for the underdamped case when we transform the co-ordinates $[u,v]^{T}$ under a suitable linear map (see (zhang2023improveddiscretizationanalysisunderdamped, , Lemma 16).
@@ 5 Proof for Overdamped PLMC
Let $X_{tK+i}$ denote the iterates of overdamped Langevin Monte-Carlo with step-size $\frac{\eta}{K}$. Let ${\tilde{X}}_{tK+i}$ denote the iterates of Poisson overdamped Langevin Monte-Carlo with step size $\eta$, and $\hat{\tilde{X}}_{tK+i}$ denote the cheap interpolants as defined in Section 2.
@@
Let Assumption 1 hold and let $\frac{\alpha\eta}{K}<1$. Let $V=\sqrt{\tfrac{\eta K}{2}}(H_{i,t}-1/K)[\nabla F(y_{0})-\nabla F(\hat{y})]$. Then we have,
$\mathbb{E}[||X_{tK+i+1}-\tilde{X}_{tK+i+1}||^{2}|\mathcal{G}]\leq(1-\tfrac{\alpha\eta}{2K})^{2}||x-y||^{2}+\tfrac{9\eta L^{2}}{\alpha K}||y-\hat{y}||^{2}+\tfrac{2\eta}{K}\mathbb{E}[||Y_{tK+i}+V-Z_{tK+i}||^{2}|\mathcal{G}].$
@@
Let $\nu=\text{Tr}(VV^{T}|\mathcal{G})$ with $V$ defined as in Proposition 1, and $\beta^{2}=\frac{\eta KL^{2}}{2}||y_{0}-\hat{y}||^{2}.$ Let $S\in\mathcal{G}^{\sigma}$ be an event. Conditioned on $\mathcal{G}$, there exists a coupling of $(Y_{tK+i},H_{i,t})$ and $Z_{tK+i}$ such that
$\mathbb{E}[||Y_{tK+i}+V-Z_{tK+i}||^{2}|\mathcal{G}]\leq(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot 2\nu+\mathbf{1}_{S^{c}}\cdot\frac{11}{2}\nu^{2}.$
@@
$E\lesssim(\eta^{6}L^{4}d+\tfrac{\eta^{4}L^{2}}{K}+\tfrac{\eta^{5}L^{4}}{\alpha K})\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}+\tfrac{\eta^{4}L^{4}d}{\alpha K}+\tfrac{\eta^{5}L^{4}d^{2}}{K}+\exp(c_{1}d-(c_{2}\eta^{2}L^{2}K)^{-1})\cdot\tfrac{\eta^{3}L^{2}d}{K},$
@@ 5.1 Finishing the proof
$W_{2}^{2}(X_{tK},\tilde{X}_{tK})$ $\lesssim\exp(-\alpha\eta t)W_{2}^{2}(X_{0},\tilde{X}_{0})^{2}+(\eta^{6}L^{4}Kd+{\eta^{4}L^{2}}+\tfrac{\eta^{5}L^{4}}{\alpha})\sum_{s=0}^{t-1}\mathbb{E}||\nabla F(\tilde{X}_{Ks})||^{2}$
@@
Note that $X_{0}=\tilde{X}_{0}$, so $W_{2}^{2}(X_{0},\tilde{X}_{0})=0.$ Moreover, $\sum_{t=0}^{N}\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}$ is bounded in Lemma 9. Substituting this above completes the proof of Theorem 1.
@@ 6 Proof for Underdamped PLMC
with $\hat{\tilde{X}}_{tK+i}^{\prime}=A_{\frac{i\eta}{K}}^{\prime}\tilde{X}_{tK}^{\prime}+G_{\frac{i\eta}{K}}^{\prime}b(\tilde{X}_{tK}^{\prime})+\sum_{j=0}^{i-1}A_{\frac{\eta(i-j-1)}{K}}^{\prime}\Gamma_{\tfrac{\eta}{K}}^{\prime}Y_{tK+j}.$ Our proof follows the same method as in the overdamped case. As before, $(Z_{tK+i})_{t,i}$ are i.i.d. $\mathcal{N}(0,\mathbf{I}_{d})$ and so are $(Y_{tK+i})_{t,i}$. ($X_{0}$, $(Z_{tK+i})_{t,i}$) and ($\tilde{X}_{0}$, $(H_{i,t})_{t,i}$$(Y_{tK+i})_{t,i}$) live on the same probability space with an arbitrary coupling. We apply will Lemma 5 to produce a particular coupling.
@@
The following Proposition is proved in Section G.1. Writing the ULMC updates as $X^{\prime}_{tK+i+1}=T(X^{\prime}_{tK+i})+\Gamma^{\prime}_{\eta/K}Z_{tK+i}$, the proof of the following Proposition hinges on the fact that $T$ is a contraction mapping (i.e, Liphschitz constant $<1$).
@@ 6.1 Finishing the proof
Note that $X_{0}^{\prime}=\tilde{X}_{0}^{\prime}$ by definition, so the first term is zero. The moments $\sum_{s=0}^{t}\mathbb{E}||\tilde{V}_{tK}||^{2p}$ and $\sum_{s=0}^{t}\mathbb{E}||\nabla F(\tilde{U}_{tK})||^{2p}$ are bounded in Theorem 6. Substitute them in, ignoring lower order terms via the assumption $\gamma\eta<c_{0}.$
@@ 7 Discussion and limitations:
Our work obtains novel convergence bounds for a randomized midpoint type method for the classical, well studied problem of strong log-concave sampling with Langevin dynamics. We improve the dependence on $\epsilon$ compared to all priors works and show that strong $L^{2}$ error lower bounds are not fundamental limits for sampling algorithms. However, the dependence on dimension $d$ and the condition number $\kappa$ in our bounds is slightly worse than prior works. We believe that further work is required to obtain better dependence on these parameters.
@@ Appendix A General Technical Lemmas and Proofs
Let $n>0$ and let $Y$ be an $\mathbb{R}^{k}$ valued random variable with mean 00, covariance $\Sigma/n$ and $\|Y\|\leq\frac{\beta}{\sqrt{n}}$ almost surely. For $t\geq 0$, let $Z_{t}$ denote a Gaussian of mean 0 and covariance $t\Sigma$ independent of $Y$. Let $\sigma_{\text{min}}^{2}$ denote the smallest eigenvalue of $\Sigma$. Then, for any $n\geq\frac{5\beta^{2}}{\sigma_{\text{min}}^{2}},$ we have
@@ A.1 Proof of Lemma 1
$2W_{2}^{2}(\sqrt{\mathbf{I}_{d}+\Sigma}W,Z)^{2}=4+2\nu-4\sqrt{1+\nu}\leq\frac{1}{2}\nu^{2}.$
@@ Appendix B Technical Results for OLMC
Let $M_{t,K}=\sup_{0\leq j<K}||\sum_{i=0}^{j}\sqrt{\frac{2\eta}{K}}Y_{tK+i}||$, and $p\in\mathbb{N}.$ Let $N_{t}:=\sum_{i=0}^{k-1}H_{i,t}.$ Then the following bounds are true.
$\sup_{0\leq i\leq k-1}||\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}||$ $\leq\eta||\nabla F(\tilde{X}_{tK})||+M_{t,k}.$
$\sup_{0\leq i\leq k-1}||\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK+i}||$ $\leq\eta LN_{t}\sup_{i\leq k-1}||\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}||.$
@@
$\frac{\eta}{K}(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot\nu+\mathbf{1}_{S^{c}}\cdot\nu^{2}\lesssim(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot\frac{\eta^{2}L^{2}}{K}\|y_{0}-\hat{y}\|^{2}+\mathbf{1}_{S^{c}}\cdot\frac{\eta^{3}L^{4}}{K}\|y_{0}-\hat{y}\|^{4}$
@@
$\eta\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim[F(\tilde{X}_{tK})-F(X_{(t+1)K})]+\eta Ld.$
@@ C.1 Proof of Proposition 1
$X_{tK+i+1}=T(X_{tK+i})+\sqrt{\frac{2\eta}{K}}Z_{tK+i},\text{ and }\tilde{X}_{tK+i+1}=T(\tilde{X}_{tK+i})+\sqrt{\frac{2\eta}{K}}\tilde{Z}_{tK+i}.$
@@ C.2 Proof of Corollary 1
Wasserstein-2 is a metric on the space of probability measures over $\mathbb{R}^{d}$ with finite second moments. By triangle inequality
$W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\pi)$ $\leq(W_{2}(\mathsf{Law}(\tilde{X}_{NK}),\mathsf{Law}(X_{NK}))+W_{2}(\mathsf{Law}(X_{NK}),\pi))^{2}$
$\lesssim W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\mathsf{Law}(X_{NK}))+W_{2}^{2}(\mathsf{Law}(X_{NK}),\pi).$
@@
By our choice of $K$, we have $\frac{\eta}{K}=\frac{\epsilon^{2}}{4L},$ and by the above Theorem, we have $W_{2}^{2}(X_{NK},\pi)\leq\frac{\epsilon^{2}d}{\alpha}$ for $N=\log(\frac{2W_{2}^{2}(X_{0},\pi)\alpha}{\epsilon^{2}d})(\alpha\eta)^{-1}.$ Let $L_{1}=2\max(C_{F},\log(\frac{2W_{2}^{2}(X_{0},\pi)\alpha}{\epsilon^{2}d})).$ Now apply Theorem 1 with
@@
and $N$ as above, to see $W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\mathsf{Law}(X_{NK}))\lesssim\frac{\epsilon^{2}d}{\alpha}.$ We assume
$\epsilon\lesssim\min\Big{(}1,\kappa^{-1/6}L_{1}^{-1/6},d^{-1/8}L_{1}^{-1/8},\kappa^{-1/4},d^{-1/4}\Big{)},$
which is a regime in which the batch size $K=\frac{4\eta L}{\epsilon^{2}}\geq 1.$
@@ C.3 Proof of Lemma 2
where $E\lesssim\frac{\eta L^{2}}{\alpha K}\mathbb{E}||{\tilde{X}}_{tK+i}-\hat{\tilde{X}}_{tK+i}||^{2}+\frac{\eta}{K}\mathbb{E}[(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot 2\nu+\mathbf{1}_{S^{c}}\cdot\frac{11}{2}\nu^{2}]$ and $S\in\mathcal{G}^{\sigma}$ is any event. We choose a particular event $S$ and bound the latter term in Lemma 7. The former term is bounded below, using items 1 and 2 of Lemma 6.
$\frac{\eta L^{2}}{\alpha K}\mathbb{E}||{\tilde{X}}_{tK+i}-\hat{\tilde{X}}_{tK+i}||^{2}\lesssim\frac{\eta^{3}L^{4}}{\alpha K}\mathbb{E}\left[N_{t}^{2}\sup_{j\leq K-1}||\tilde{X}_{tK}-\hat{\tilde{X}}_{tK+j}||^{2}\right]$
@@
$\frac{\eta L^{2}}{\alpha K}\mathbb{E}||{\tilde{X}}_{tK+i}-\hat{\tilde{X}}_{tK+i}||^{2}$ $\lesssim\frac{\eta^{5}L^{4}}{\alpha K}\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}+\frac{\eta^{3}L^{4}}{\alpha K}\mathbb{E}[M_{t,K}^{2}]$
$\lesssim\frac{\eta^{5}L^{4}}{\alpha K}\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}+\frac{\eta^{4}L^{4}d}{\alpha K}.$
@@ Appendix D Technical Results for ULMC
[…]
@@ Appendix F Proof of Corollary 2
The assumption $\epsilon\lesssim\min(1,\kappa d^{-1/2},\kappa^{\frac{p+3}{2(3p+1)}}d^{-\frac{p}{3p+1}})$ ensures that the batch size $K\asymp\frac{\eta L}{\epsilon\sqrt{\alpha}}\geq 1$ making our algorithm valid. Since $W_{2}$ is a metric, triangle inequality gives
@@
With $K$ defined as in the corollary we have $\frac{\eta}{K}=\frac{0.94\epsilon\sqrt{\alpha}}{L\sqrt{2}},$ thus $W_{2}^{2}(U_{NK},\pi)\leq\frac{\epsilon^{2}d}{\alpha}$ for $N\geq\frac{c_{3}\gamma}{\alpha h}\log\frac{\sqrt{2\alpha}W_{2}(\nu_{0},\pi)}{\epsilon\sqrt{d}}.$ It remains to be shown that $W_{2}^{2}(\tilde{X}_{NK},X_{NK})\lesssim\frac{\epsilon^{2}d}{\alpha}.$ Let $k$ be a natural number. Under our assumptions on $V_{0}$ and $X_{0}$, we have $\mathbb{E}||V_{0}||^{2k}=d^{k}$ and
@@
$\leq L||\Psi_{0}-x^{*}||^{2}$
$\lesssim L||X_{0}-x^{*}||^{2}+\frac{L}{\gamma^{2}}||V_{0}||^{2}$
Under our assumptions, we thus get $\mathbb{E}|(F(\Psi_{0})-F(\Psi_{T}))^{+}|^{k}\lesssim d^{k}.$ Moreover, we have $\log K=\log\frac{\eta L}{\epsilon\sqrt{\alpha}}\lesssim\log\frac{\kappa}{\epsilon}$ under the assumption that $\gamma\eta<c_{0}.$ Now let $L_{2}=c_{2}\log\frac{\sqrt{2\alpha}W_{2}(\nu_{0},\pi)}{\epsilon\sqrt{d}},$ $L_{3}=\log\frac{\kappa}{\epsilon}$ and apply Theorem 3 with $N$ as above, and
@@ G.1 Proof of Proposition 2
[…]

### v3

@@ 4 Proof overview
    $V||\leq\beta$ a.s., $\mathbb{E}[V]=0$, and $\mathbb{E}[VV^{T}]=\Sigma.$
@@
[…]
@@
The proof in this case is similar to the overdamped case. Here we use the fact we can obtain a contraction map similar to $T$ in Equation (4) even for the underdamped case when we transform the co-ordinates $u,v]^{T}$ under a suitable linear map (see (zhang2023improveddiscretizationanalysisunderdamped, , Lemma 16).
@@ 5 Proof for Overdamped PLMC
Let $X_{tK+i}$ denote the iterates of overdamped Langevin Monte Carlo with step-size $\frac{\eta}{K}$. Let ${\tilde{X}}_{tK+i}$ denote the iterates of Poisson overdamped Langevin Monte Carlo with step size $\eta$, and $\hat{\tilde{X}}_{tK+i}$ denote the cheap interpolants as defined in Section 2.
@@
[…]
@@
[…]
@@
[…]
@@ 5.1 Finishing the proof
[…]
@@
[…]
@@ 6 Proof for Underdamped PLMC
with $\hat{\tilde{X}}_{tK+i}^{\prime}=A_{\frac{i\eta}{K}}^{\prime}\tilde{X}_{tK}^{\prime}+G_{\frac{i\eta}{K}}^{\prime}b(\tilde{X}_{tK}^{\prime})+\sum_{j=0}^{i-1}A_{\frac{\eta(i-j-1)}{K}}^{\prime}\Gamma_{\tfrac{\eta}{K}}^{\prime}Y_{tK+j}.$ Our proof follows the same method as in the overdamped case. As before, $(Z_{tK+i})_{t,i}$ are i.i.d. $\mathcal{N}(0,\mathbf{I}_{d})$ and so are $(Y_{tK+i})_{t,i}$. ($X_{0}$, $(Z_{tK+i})_{t,i}$) and ($\tilde{X}_{0}$, $(H_{i,t})_{t,i}$$(Y_{tK+i})_{t,i}$) live on the same probability space with an arbitrary coupling. We will apply Lemma 1 to produce a particular coupling.
@@
The following Proposition is proved in Section G.1. Writing the ULMC updates as $X^{\prime}_{tK+i+1}=T(X^{\prime}_{tK+i})+\Gamma^{\prime}_{\eta/K}Z_{tK+i}$, the proof hinges on the fact that $T$ is a contraction mapping (i.e, Liphschitz constant <1<1).
@@ 6.1 Finishing the proof
[…]
@@ 7 Discussion and limitations:
Our work obtains novel convergence bounds for a randomized midpoint type method for the classical, well studied problem of strong log-concave sampling with Langevin dynamics. We improve the dependence on $\epsilon$ compared to all prior works and show that strong $L^{2}$ error lower bounds are not fundamental limits for sampling algorithms. However, the dependence on dimension $d$ and the condition number $\kappa$ in our bounds is slightly worse than prior works. We believe that further work is required to obtain better dependence on these parameters.
@@ Appendix A General Technical Lemmas and Proofs
Let $n>0$ and let $Y$ be an $\mathbb{R}^{k}$ valued random variable with mean 0, covariance $\Sigma/n$ and $\|Y\|\leq\frac{\beta}{\sqrt{n}}$ almost surely. For $t\geq 0$, let $Z_{t}$ denote a Gaussian of mean 0 and covariance $t\Sigma$ independent of $Y$. Let $\sigma_{\text{min}}^{2}$ denote the smallest eigenvalue of $\Sigma$. Then, for any $n\geq\frac{5\beta^{2}}{\sigma_{\text{min}}^{2}},$ we have
@@ A.1 Proof of Lemma 1
$2W_{2}^{2}(\sqrt{\mathbf{I}_{d}+\Sigma}W,Z)=4+2\nu-4\sqrt{1+\nu}\leq\frac{1}{2}\nu^{2}.$
@@ Appendix B Technical Results for OLMC
[…]
@@
$\frac{\eta}{K}[(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot\nu+\mathbf{1}_{S^{c}}\cdot\nu^{2}]\lesssim(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot\frac{\eta^{2}L^{2}}{K}\|y_{0}-\hat{y}\|^{2}+\mathbf{1}_{S^{c}}\cdot\frac{\eta^{3}L^{4}}{K}\|y_{0}-\hat{y}\|^{4}$
@@
$\eta\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim[F(\tilde{X}_{tK})-F(\tilde{X}_{(t+1)K})]+\eta Ld.$
@@ C.1 Proof of Proposition 1
[…]
@@ C.2 Proof of Corollary 1
Wasserstein-2 is a metric on the space of probability measures over $\mathbb{R}^{d}$ with finite second moments. By triangle inequality,
$W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\pi)\lesssim W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\mathsf{Law}(X_{NK}))+W_{2}^{2}(\mathsf{Law}(X_{NK}),\pi).$
@@
By our choice of $K$, we have $\frac{\eta}{K}\lesssim\frac{\epsilon^{2}}{L}.$ Note that the above Theorem goes through with an inequality $\eta\leq\frac{\epsilon^{2}}{4L}$, so we have $W_{2}^{2}(X_{NK},\pi)\leq\frac{\epsilon^{2}d}{\alpha}$ for $N=\log(\frac{2W_{2}^{2}(X_{0},\pi)\alpha}{\epsilon^{2}d})(\alpha\eta)^{-1}.$ Let $L_{1}=2\max(C_{F},\log(\frac{2W_{2}^{2}(X_{0},\pi)\alpha}{\epsilon^{2}d})).$ Now apply Theorem 1 with
@@
and $N$ as above, to see $W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\mathsf{Law}(X_{NK}))\lesssim\frac{\epsilon^{2}d}{\alpha}.$
@@ C.3 Proof of Lemma 2
[…]
@@
[…]
@@ Appendix D Technical Results for ULMC
Let $\Pi:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ denote projection onto the first $d$ coordinates. Let $G_{h}$ and $A_{h}$ be as defined in the update rule for underdamped Langevin Monte-Carlo in Section 2. Let $\hat{\tilde{X}}_{tK+i}$ and ${\tilde{X}}_{tK+i}$ denote the cheap interpolation and iterates respectively of Poisson-ULMC, as defined in Section 2. Let $\|\cdot\|$ denote the operator norm of a matrix, and $\|\cdot\|_{\mathbb{R}^{n}}$ denote the Euclidean norm in dimension $n$. Then the following inequalities are true.
@@ Appendix F Proof of Corollary 2
Since $W_{2}$ is a metric, triangle inequality gives
@@
[…]
@@
[…]
Under our assumptions, we thus get $\mathbb{E}|(F(\Psi_{0})-F(\Psi_{T}))^{+}|^{k}\lesssim d^{k}.$ Moreover, we have $\log K\asymp\max(0,\log\frac{\eta L}{\epsilon\sqrt{\alpha}})\lesssim\log\frac{\kappa}{\epsilon}$ under the assumption that $\gamma\eta<c_{0}.$ Now let $L_{2}=c_{2}\log\frac{\sqrt{2\alpha}W_{2}(\nu_{0},\pi)}{\epsilon\sqrt{d}},$ $L_{3}=\log\frac{\kappa}{\epsilon}$ and apply Theorem 3 with $N$ as above, and
@@ G.1 Proof of Proposition 2
Let $\|\cdot\|$ denote the operator norm of a matrix, and $\|\cdot\|_{\mathbb{R}^{n}}$ denote the Euclidean norm in dimension $n$. Let $p=[p_{u},p_{v}]^{T}$ and $q=[q_{u},q_{v}]^{T}$ denote arbitrary points in $\mathbb{R}^{2d}$. Assume $\gamma h<c_{0}$ for some sufficiently small constant $c_{0}>0,$ and Assumption 1. Then the following inequalities are true.